mcq A boiler which produces water vapour at atmospheric pressure, has a 2 cm thick ,...

mcq

A boiler which produces water vapour at atmospheric pressure, has a 2 cm thick , 500 cm2bottom, which is made up of a metal whose K= 400 Wm-1 0C-1. A heating coil of 20 kW supplies heat to the boiler. If no heat is lost to the outside and the latent heat of vaporisation L=2.3 x 106 Jkg-1, find the maximum rate of production of water vapour in kgs-1?

(1) 3.07 x 10-1 (2) 8.7 x 10-3 (3) 4.2 x 10-2 (4) 8.7 x 10-4 (5) 2.3 10-3

need a detailed answer
Answer #1

See for maximum water vapour production the condition is that the

Total heat produced by coil= heat absorbed by water

If water vapour production rate is m then

m× heat of vaporization= heat produced by coil

m×2.3×106= 20×103. (1 KW= 1000 W)

m=8.7×10-3 kg/s

Answer is option 2.